# Stochastic Subspace via Probabilistic Principal Component Analysis for Characterizing Model Error

### v1

@@ Abstract
This paper proposes a probabilistic model of subspaces based on the probabilistic principal component analysis (PCA). Given a sample of vectors in the embedding space—commonly known as a snapshot matrix—this method uses quantities derived from the probabilistic PCA to construct distributions of the sample matrix, as well as the principal subspaces. It is applicable to projection-based reduced-order modeling methods, such as proper orthogonal decomposition and related model reduction methods. The stochastic subspace thus constructed can be used, for example, to characterize model-form uncertainty in computational mechanics. The proposed method has multiple desirable properties: (1) it is naturally justified by the probabilistic PCA and has analytic forms for the induced random matrix models; (2) it satisfies linear constraints, such as boundary conditions of all kinds, by default; (3) it has only one hyperparameter, which significantly simplifies training; and (4) its algorithm is very easy to implement. We compare the proposed method with existing approaches in a low-dimensional visualization example and a parametric static problem, and demonstrate its performance in a dynamics model of a space structure.
@@ 1 Introduction
Internal direct representation methods improve a model through intrusive modifications, rather than by adding an extra term to the output. Embedded model enrichment [14] provides a framework for extrapolative predictions using composite models. Morrison et al. [16] address model inadequacy with stochastic operators introduced as a source in the governing equations while preserving underlying physical constraints. Similarly, Portone and Moser [17] propose a stochastic operator for an unclosed dispersion term in an averaged advection–diffusion equation for model closure. However, modifying the governing equations of the model brings challenges in the computational cost and implementation. Additionally, significant prior knowledge of the system is required to alter the governing equations and achieve improvements effectively. Bhat et al. [18] address model-form error by embedding dynamic discrepancies within small-scale models, using Bayesian smoothing splines (BS-ANOVA) framework. Strong and Oakley [11] introduce a novel way to quantify model inadequacies by decomposing the model into sub-functions and addressing discrepancies in each sub-function. The authors show the application of their method in health economics studies. Sargsyan et al. [19] improves uncertainty representation and quantification in computational models by introducing a Bayesian framework with embedded model error representation using polynomial chaos expansion. Other studies have used internal model error correction for various application areas, e.g., Reynolds-averaged Navier Stokes (RANS) applications [20], large eddy simulation (LES) computations [21], molecular simulations [22], particle-laden flows [23], and chemical ignition modeling [24]. In general, internal direct representation methods require a deep understanding of the system to implement the changes necessary for improvement. This knowledge is not always available and often demands an extensive study of the system.
@@
A recent work by Soize and Farhat [32] in model-form uncertainty merges ideas from projection-based reduced-order modeling [33, 34] and random matrix theory [35, 36]. In the new approach, instead of randomizing coefficient matrices, the reduced-order basis (ROB) is randomized. The key observation is that since the ROB determines the ROM, randomizing the ROB also randomizes the ROM, which can be used for efficient probabilistic analysis of model error. Using computationally inexpensive ROM instead of a full-order model makes the approach computationally tractable for uncertainty quantification tasks. In this context, a few probabilistic models and estimation procedures have been proposed to construct stochastic ROBs [32, 37], and applied to various engineering problems [38, 39, 40]. However, they may face limitations due to high numbers of hyperparameters, complex implementation, or poor uncertainty quantification (UQ). Furthermore, the previous studies [32, 37] have some hyperparameters tuned by the trial-and-error method, which can be inaccurate and computationally expensive. This study aims to develop a stochastic subspace model that is free from the aforementioned limitations.
@@
    We use this stochastic subspace model to construct stochastic ROMs, and reduce the number of hyperparameters to a single scalar. This is in contrast to existing stochastic ROM approaches, which use stochastic ROBs, and often suffer from complex implementation and large numbers of hyperparameters.
@@
    The optimization of the hyperparameter is fully automated and done efficiently by exploiting one-dimensional optimization algorithms. Existing approaches, however, can have hyperparameter tuning procedures that are computationally intensive and often have some hyperparameters that are manually tuned.
@@
    We demonstrate in a series of numerical experiments that our method provides consistent and significantly sharper uncertainty estimates, with much lower computational costs.
This paper is organized as follows: A brief introduction to stochastic reduced-order modeling is presented in Section 2. The proposed stochastic subspace model, along with the algorithm, is described in Section 3. Critical information regarding the hyperparameter optimization is given in Section 4. Comparisons with related works in the literature are done in Section 5. The accuracy and efficiency of the proposed method are validated using numerical examples in Section 6. Finally, Section 7 concludes the paper with a brief summary and potential future work.
@@ 4 Hyperparameter training
where $\mathbf{u}_{\_}E$ is the experimental or ground-truth observation of the output, $\mathbf{u}_{\_}L$ is the low-fidelity prediction of the SROM, and $d_{\_}o(\mathbf{u}):=\|\mathbf{u}-\mathbf{u}_{\_}L^{o}\|_{\_}{L^{2}}$ is the $L^{2}$ distance to the low-fidelity prediction $\mathbf{u}_{\_}L^{o}$ of a reference model. This objective function aims to improve the consistency of the SROM in characterizing the error of the reference model.
@@
5 Related work and comparison
In this section, the two main works that use SROM for model error characterization are discussed in detail. Firstly, the work by Soize and Farhat [32] is discussed in Section 5.1. In Section 5.2, the work by Zhang and Guilleminot [37] to characterize model-form uncertainties across multiple models is presented. In Section 5.3, we distinguish the three methods based on their operational mechanisms and their dependence on available information. Then, in Section 5.4, the two main metrics used to compare the performance of different methods are described.
@@ 5.1 Non-parametric model (NPM)
The stochastic basis in the NPM [32] can be written as:
@@
The random matrix $\mathbf{G}$ requires knowledge of the discretization $\mathscr{D}$ of the spatial fields, and is sampled from a Gaussian process with a hyperparameter $\gamma$ for correlation length-scale. The column dependence matrix $\mathbf{R}$ is an order-$k$ upper triangular matrix with positive diagonal entries. The scale $s\in[0,\infty)$ is a non-negative number. Once $\mathbf{U}$ is constructed, orthogonal projection $\mathbf{P}_{\_}{\mathscr{B}^{\perp}}$ enforces the linear constraints, tangential projection $\mathcal{P}^{\text{St}}_{\_}{\overline{\mathbf{V}}}$ projects a matrix to the tangent space of the Stiefel manifold at a reference ROB $\overline{\mathbf{V}}$, and orthonormalization $\pi$ as in Eq. 4 gives an orthonormal basis. The hyperparameters $(s,\gamma,\mathbf{R})$ are trained by minimizing a weighted average of $J_{\_}{\text{mean}}$ and $J_{\_}{\text{std}}$: the former aims at matching the SROM mean prediction to ground-truth observations $\mathbf{O}$; the latter aims at matching the SROM standard deviation to a scaled difference between $\mathbf{O}$ and $\mathbf{u}_{\_}L^{o}$.
The seminal work of NPM introduced the idea of using SROM to analyze model-form uncertainty, but as we will illustrate in Section 6, it has several challenges. Its number of hyperparameters is $k(k+1)/2+2$—which can be quite large—and its objective function has many local optima; both lead to significant difficulty in optimization. Its sampling requires knowledge of spatial discretization, which can be hard to obtain in implementation. The concentration of its distribution is controlled by the scale and the weight in the objective function, which are manually tuned and difficult to choose. Moreover, the linear constraints are enforced independently of the physical model, which may lead to physically incoherent SROM predictions.
@@ 5.2 Riemannian stochastic model (RSM)
To characterize model-form uncertainty given multiple plausible models, the RSM [37] uses the following stochastic basis:
@@
where $c\in[0,\infty)$ is a scale parameter, $q$ is the number of models, $\mathbf{p}=(p_{\_}i)_{\_}{i=1}^{q}$ is a probability vector distributed as $\text{Dirichlet}(\boldsymbol{\alpha})$ with concentration parameters $\boldsymbol{\alpha}\in\mathbb{R}_{\_}{>0}^{q}$. The Riemannian logarithm $\log^{\text{St}}_{\_}{\overline{\mathbf{V}}}$ maps a model-specific ROB $\mathbf{V}^{(i)}$ to the tangent space of the Stiefel manifold at a reference ROB $\overline{\mathbf{V}}$. The Riemannian exponential $\exp^{\text{St}}_{\_}{\overline{\mathbf{V}}}$ maps a tangent vector at $\overline{\mathbf{V}}$ back to the Stiefel manifold, giving the orthonormal basis $\mathbf{W}$. The hyperparameters $\boldsymbol{\alpha}$ are trained by minimizing $\|\mathbb{E}[\log^{\text{St}}_{\_}{\overline{\mathbf{V}}}\mathbf{W}]\|_{\_}F^{2}$, to match the center of mass of the distribution to $\overline{\mathbf{V}}$. It is a quadratic program and can be solved efficiently. The scale parameter $c$, on the other hand, is tuned manually.
The RSM is significantly simpler than the NPM, and its number of hyperparameters—$q+1$—is often small, which makes training easier. However, it is not applicable outside multi-model settings, and it may not give consistent UQ.
@@ 5.3 Comparison of model structures
Figure 1 outlines the structures of the three methods. Both NPM and RSM construct probabilistic models of reduced-order bases, while SS-PPCA models subspaces. The NPM requires discretization information and boundary conditions for hyperparameter optimization and SROM sampling, while SS-PPCA and RSM do not. The RSM optimizes its hyperparameters using information only from the reduced bases, but SS-PPCA and NPM also use observed quantities of interest to define their objective functions.
5.4 Metrics for comparing stochastic models
Two general metrics for quantifying model-form uncertainty are consistency and sharpness. By consistent UQ, we mean that the statistics on prediction uncertainty—such as standard deviation and predictive intervals (PI)—derived from the model match the prediction error on average. Without consistency, the model either under- or over-estimates its prediction error, which risks being over-confident or unnecessarily conservative. By sharp UQ, we mean the model’s PIs are narrower than alternative methods. This is essentially the probabilistic version of model accuracy and is desirable as we want to minimize error.
@@ 6 Numerical experiments
In this section, we compare the proposed method, SS-PPCA, with the already available methods, NPM and RSM, using four numerical examples. The examples are in increasing order of complexity. We begin with a low-dimensional illustrative example presented in Section 6.1. In Section 6.2, we examine a parametric linear static problem and compare all three methods. In Section 6.3, we demonstrate how our approach can be used to characterize the HDM error, assuming experimental data is available. Finally, in Section 6.4, we discuss a linear dynamics problem involving a space structure. All the data and codes for these examples are available at https://github.com/UQUH/SSPPCA.
6.1 Illustrative example: sampling on $\text{St}(3,1)$
To gain a clearer understanding of the differences among SS-PPCA, NPM, and RSM, we design a simple, low-dimensional example, where the reduced bases are on $\text{St}(3,1)$—the unit sphere. This example aids in visualizing the stochastic reduced-order bases (SROBs) or subspaces generated by each method, which is not possible in other examples due to the high dimensionality.
Multiple data-generating models are used so that we can compare the performance of all three methods. The population models are given by $\mathbf{x}^{(j)}\sim N_{\_}3(0,\boldsymbol{\Sigma}^{(j)})$, $j=1,\cdots,q$. Covariance matrices $\boldsymbol{\Sigma}^{(j)}$ are sampled from the Wishart distribution $W_{\_}3(\operatorname{diag}(\boldsymbol{\sigma})^{2},p)$ as follows:
  1. 1.
    Input: degree of freedom $p\in\{3,4,\cdots\}$ and standard deviation vector $\boldsymbol{\sigma}\in\mathbb{R}_{\_}{>0}^{3}$.
  2. 2.
[…]
  3. 3.
    $\mathbf{Y}^{(j)}\leftarrow\operatorname{diag}(\boldsymbol{\sigma})\,\mathbf{Z}^{(j)}$.
  4. 4.
    $\boldsymbol{\Sigma}^{(j)}\leftarrow\mathbf{Y}^{(j)}{\mathbf{Y}^{(j)}}^{\intercal}$.
  5. 5.
    Output: a random matrix $\boldsymbol{\Sigma}^{(j)}\sim W_{\_}3(\operatorname{diag}(\boldsymbol{\sigma})^{2},p)$.
[…]
The ROM data is derived by projecting the HDM data onto a one-dimensional subspace:
$\mathbf{X}_{\_}R=\overline{\mathbf{v}}\,\overline{\mathbf{v}}^{\intercal}\mathbf{X},$ (11)
where $\overline{\mathbf{v}}=\mathbf{v}_{\_}1(\mathbf{X})$ is the global POD basis. Here, $\mathbf{v}_{\_}1(\mathbf{M})$ denotes the first principal component associated with the largest singular value of a matrix $\mathbf{M}$. It can be computed either via a truncated SVD of $\mathbf{M}$ or a truncated symmetric EVD of $\mathbf{M}\mathbf{M}^{\intercal}$; the latter can be more accurate and efficient when $\mathbf{M}$ is wide.
The SROBs for the NPM method are generated using the strategy described in [32], where $\mathbf{G}\sim N_{\_}3(0,\mathbf{I}_{\_}3)$, $\mathbf{U}=s\mathbf{G}$ and $\mathbf{B}=0$. For the RSM method, the SROBs are generated using the strategy described in [37], with model-specific basis $\mathbf{v}^{(j)}=\mathbf{v}_{\_}1(\mathbf{X}^{(j)})$, $j=1,\cdots,q$. The stochastic subspaces for our SS-PPCA method are generated using the stochastic subspace model described in Section 3.3. The hyperparameter $\beta$ of our method is optimized efficiently using the strategy given in Section 4, which is found to be five.
In this example, ROBs and subspaces can be represented as points on the unit sphere. For the visualization, vectors on the unit sphere have their signs flipped to be on the hemisphere centered around $\overline{\mathbf{v}}$, and projected on to the $x_{\_}1x_{\_}2$-plane. The unit sphere is defined as $\mathbb{S}^{2}:=\{\mathbf{x}\in\mathbb{R}^{3}:|\mathbf{x}|=1\}$. The closed hemisphere _$\mathbb{H}^{2}_{\_}\mathbf{p}$ centered around a point $\mathbf{p}\in\mathbb{S}^{2}$ is defined as _$\mathbb{H}^{2}_{\_}\mathbf{p}:=\{\mathbf{x}\in\mathbb{S}^{2}:\langle\mathbf{x},\mathbf{p}\rangle\geq 0\}$. Projection $P:\mathbb{S}^{2}\mapsto\mathbb{H}^{2}_{\_}\mathbf{p}$ is defined as: $P(\mathbf{x})=\mathbf{x}$ if $\langle\mathbf{x},\mathbf{p}\rangle\geq 0$, and $P(\mathbf{x})=-\mathbf{x}$ if $\langle\mathbf{x},\mathbf{p}\rangle<0$. Figure 3 compares stochastic ROBs of the three methods. We observe that the NPM samples are isotropic even though the HDM data itself is anisotropic. The RSM samples are more concentrated and anisotropic since they are, by default, contained within the convex hull of the model-specific ROBs, which are anisotropically distributed. The RSM method can generate samples outside the convex hull, but the scaling hyperparameter $c$ needs to be tuned manually. Hence, we restrict $c$ = 1 to generate the RSM sample. The SS-PPCA samples are anisotropic and have the highest spread among the methods.
Figure 2: SROB samples from the three methods.
Figure 3: Probability density functions of the $L^{2}$ distance to ROM prediction.
For consistency, ideally the SROM should match the HDM so that the latter can be seen as a random sample of the former. However, since the ROM dimension is smaller than the HDM dimension ($k\ll n$), the SROM can never capture the full complexities of the HDM. Instead, we compare the behaviors of HDM and SROM in terms of quantities of interest. Figure 3 compares distributions of the $L^{2}$ distance of SROM predictions to the global ROM predictions. If a SROM is consistent, HDM should appear similar to a typical SROM. However, only SS-PPCA passes this hypothesis test: the tail probabilities estimated using $10^{6}$ samples are SS-PPCA 10.4%, NPM 1.35%, and RSM 0.0%.

### v2

@@ Abstract
This paper proposes a probabilistic model of subspaces based on the probabilistic principal component analysis (PCA). Given a sample of vectors in the embedding space—commonly known as a snapshot matrix—this method uses quantities derived from the probabilistic PCA to construct distributions of the sample matrix, as well as the principal subspaces. It is applicable to projection-based reduced-order modeling methods, such as proper orthogonal decomposition and related model reduction methods. The stochastic subspace thus constructed can be used, for example, to characterize model-form uncertainty in computational mechanics. The proposed method has multiple desirable properties: (1) it is naturally justified by the probabilistic PCA and has analytic forms for the induced random matrix models; (2) it satisfies linear constraints, such as boundary conditions of all kinds, by default; (3) it has only one hyperparameter, which significantly simplifies training; and (4) its algorithm is very easy to implement. We demonstrate the performance of the proposed method via several numerical examples in computational mechanics and structural dynamics.
@@ 1 Introduction
Internal direct representation methods improve a model through intrusive modifications, rather than by adding an extra term to the output. Embedded model enrichment [14] provides a framework for extrapolative predictions using composite models. Morrison et al. [16] address model inadequacy with stochastic operators introduced as a source in the governing equations while preserving underlying physical constraints. Similarly, Portone and Moser [17] propose a stochastic operator for an unclosed dispersion term in an averaged advection–diffusion equation for model closure. However, modifying the governing equations of the model brings challenges in the computational cost and implementation. Additionally, significant prior knowledge of the system is required to alter the governing equations and achieve improvements effectively. Bhat et al. [18] address model-form error by embedding dynamic discrepancies within small-scale models, using Bayesian smoothing splines (BS-ANOVA) framework. Strong and Oakley [11] introduce a novel way to quantify model inadequacies by decomposing the model into sub-functions and addressing discrepancies in each sub-function. The authors show the application of their method in health economics studies. Sargsyan et al. [19] improve uncertainty representation and quantification in computational models by introducing a Bayesian framework with embedded model error representation using polynomial chaos expansion. Other studies have used internal model error correction for various application areas, e.g., Reynolds-averaged Navier Stokes (RANS) applications [20], large eddy simulation (LES) computations [21], molecular simulations [22], particle-laden flows [23], and chemical ignition modeling [24]. In general, internal direct representation methods require a deep understanding of the system to implement the changes necessary for improvement. This knowledge is not always available and often demands an extensive study of the system.
@@
A recent work by Soize and Farhat [32] in model-form uncertainty merges ideas from projection-based reduced-order modeling [33, 34] and random matrix theory [35, 36]. In the new approach, instead of randomizing coefficient matrices, the reduced-order basis (ROB) is randomized. The key observation is that since the ROB determines the ROM, randomizing the ROB also randomizes the ROM, which can be used for efficient probabilistic analysis of model error. Using computationally inexpensive ROM instead of a full-order model makes the approach computationally tractable for uncertainty quantification tasks. The probabilistic model and estimation procedure proposed in [32] have been applied to various engineering problems [37]. Because this method can face challenges due to a large number of hyperparameters and complex implementation, later developments have aimed at reducing the number of hyperparameters and simplifying hyperparameter training [38]. In settings with multiple candidate models, Zhang and Guilleminot [39] proposed a different probabilistic model for ROBs, and has been used for various problems in computational mechanics [40, 41].
@@
    We use this stochastic subspace model to construct stochastic ROMs, and reduce the number of hyperparameters to a single scalar. This contrasts with existing stochastic ROM approaches, which use stochastic bases and typically involve more hyperparameters.
@@
    The optimization of the hyperparameter is fully automated and done efficiently by exploiting one-dimensional optimization algorithms.
@@
    We demonstrate in a series of numerical experiments that our method provides consistent and sharp uncertainty estimates, with low computational costs.
This paper is organized as follows: A brief introduction to stochastic reduced-order modeling is presented in Section 2. The proposed stochastic subspace model, along with the algorithm, is described in Section 3. Critical information regarding the hyperparameter optimization is given in Section 4. Related works in the literature are reviewed in Section 5. The accuracy and efficiency of the proposed method are validated using numerical examples in Section 6. Finally, Section 7 concludes the paper with a brief summary and potential future work.
@@ 4 Hyperparameter training
where $\mathbf{u}_{\_}E$ is the experimental or ground-truth observation of the output, $\mathbf{u}_{\_}L$ is the low-fidelity prediction of the SROM, and $d_{\_}o(\mathbf{u}):=\|\mathbf{u}-\mathbf{u}_{\_}L^{o}\|_{\_}{L^{2}}$ is the $L^{2}$ distance to the low-fidelity prediction $\mathbf{u}_{\_}L^{o}$ of a reference model. Given the value of the hyperparameter $\beta$, the stochastic subspace model determines the stochastic ROM, which produces stochastic predictions $\mathbf{u}_{\_}L$ that is then summarized into the random variable $d_{\_}o(\mathbf{u}_{\_}L)$. Given the experimental or ground-truth observation $\mathbf{u}_{\_}E$, the objective function measures the mean squared error between $d_{\_}o(\mathbf{u}_{\_}L)$ and $d_{\_}o(\mathbf{u}_{\_}E)$, which is a statistical measure of how closely the SROM resembles the ground truth. Overall, this optimization problem aims to improve the consistency of the SROM in characterizing the error of the reference model.
@@
5 Related works
This section reviews in detail the two main works that use SROM for model error characterization. Firstly, the work by Soize and Farhat [32] is discussed in Section 5.1. In Section 5.2, the work by Zhang and Guilleminot [39] for characterizing model-form uncertainty across multiple models is presented. In Section 5.3, we distinguish the three methods based on their operational mechanisms and their dependence on available information.
@@ 5.1 Non-parametric model (NPM)
The seminal work of NPM [32] introduced the idea of using SROM to analyze model-form uncertainty. The stochastic basis in the NPM can be written as:
@@
The random matrix $\mathbf{G}$ requires knowledge of the discretization $\mathscr{D}$ of the spatial fields, and is sampled from a Gaussian process with a hyperparameter $\gamma$ for correlation length-scale. The column dependence matrix $\mathbf{R}$ is an order-$k$ upper triangular matrix with positive diagonal entries. The scale $s\in[0,\infty)$ is a non-negative number. Once $\mathbf{U}$ is constructed, orthogonal projection $\mathbf{P}_{\_}{\mathscr{B}^{\perp}}$ enforces the linear constraints, tangential projection $\mathcal{P}^{\text{St}}_{\_}{\overline{\mathbf{V}}}$ projects a matrix to the tangent space of the Stiefel manifold at a reference ROB $\overline{\mathbf{V}}$, and orthonormalization $\pi$ as in Eq. 4 gives an orthonormal basis. The hyperparameters $(s,\gamma,\mathbf{R})$ have a dimension of $k(k+1)/2+2$, and are trained by minimizing a weighted average of $J_{\_}{\text{mean}}$ and $J_{\_}{\text{std}}$: the former aims at matching the SROM mean prediction to ground-truth observations $\mathbf{O}$; the latter aims at matching the SROM standard deviation to a scaled difference between $\mathbf{O}$ and $\mathbf{u}_{\_}L^{o}$.
@@ 5.2 Riemannian stochastic model (RSM)
For problems where there exist a number of physically plausible candidate models, the RSM [39] characterizes model-form uncertainty using SROMs constructed with the following stochastic basis:
@@
where $c\in[0,\infty)$ is a scale parameter, $q$ is the number of models, and $\mathbf{p}=(p_{\_}i)_{\_}{i=1}^{q}$ is a probability vector following the Dirichlet distribution with concentration parameters $\boldsymbol{\alpha}\in\mathbb{R}_{\_}{>0}^{q}$. The Riemannian logarithm $\log^{\text{St}}_{\_}{\overline{\mathbf{V}}}$ maps a model-specific ROB $\mathbf{V}^{(i)}$ to the tangent space of the Stiefel manifold at a reference ROB $\overline{\mathbf{V}}$. The Riemannian exponential $\exp^{\text{St}}_{\_}{\overline{\mathbf{V}}}$ maps a tangent vector at $\overline{\mathbf{V}}$ back to the Stiefel manifold, giving the orthonormal basis $\mathbf{W}$. The scale parameter $c$ is usually set to one or omitted in later works [40, 41]. The hyperparameters $\boldsymbol{\alpha}$ are trained by minimizing $\|\mathbb{E}[\log^{\text{St}}_{\_}{\overline{\mathbf{V}}}\mathbf{W}]\|_{\_}F^{2}$, to match the center of mass of the distribution to $\overline{\mathbf{V}}$. It is a quadratic program and can be solved efficiently. We note that the RSM cannot be applied outside multi-model settings.
@@ 5.3 Comparison of model structures
To deepen understanding of the three methods, we outline their structures in Figure 1 and highlight a few key differences below. Both NPM and RSM construct probabilistic models of reduced-order bases, whereas SS-PPCA models subspaces directly. Another distinction lies in the reliance on discretization information: NPM requires discretization and boundary condition data for hyperparameter optimization and SROM sampling, while SS-PPCA and RSM do not. Finally, while RSM optimizes its hyperparameters using only the reduced bases, SS-PPCA and NPM incorporate observed quantities of interest into their objective functions.
@@ 6 Numerical experiments
In this section, we evaluate the proposed method, SS-PPCA, in characterizing model error using three numerical examples. In Section 6.1, we examine a parametric linear static problem and characterize model error at unseen parameters. In Section 6.2, we demonstrate how our approach can be used to characterize the HDM error, assuming that experimental data is available. Finally, in Section 6.3, we discuss a linear dynamics problem involving a space structure. All data and code for these examples are available at https://github.com/UQUH/SSPPCA.
The performance of the SROMs is assessed using two general metrics for stochastic models: consistency and sharpness. By consistent UQ, we mean that the statistics on prediction uncertainty—such as standard deviation and predictive intervals (PI)—derived from the model match the prediction error on average. Without consistency, the model either under- or over-estimates its prediction error, which risks being over-confident or unnecessarily conservative. By sharp UQ, we mean the model’s PIs are narrower than alternative methods. This is essentially the probabilistic version of model accuracy and is desirable as we want to minimize error.